### v2

# APTx Neuron: A Unified Trainable Neuron Architecture Integrating Activation and Computation

Abstract
We propose the APTx Neuron, a novel, unified neural computation unit that integrates non-linear activation and linear transformation into a single trainable expression. The APTx Neuron is derived from the APTx activation function, thereby eliminating the need for separate activation layers and making the architecture both computationally efficient and elegant. The proposed neuron follows the functional form $y=\sum_{i=1}^{n}((\alpha_{i}+\tanh(\beta_{i}x_{i}))\cdot\gamma_{i}x_{i})+\delta$, where all parameters $\alpha_{i}$, $\beta_{i}$, $\gamma_{i}$, and $\delta$ are trainable. We validate our APTx Neuron-based architecture on the MNIST dataset, achieving up to 96.69% test accuracy in just 20 epochs using approximately 332K trainable parameters. The results highlight the superior expressiveness and computational efficiency of the APTx Neuron compared to traditional neurons, pointing toward a new paradigm in unified neuron design and the architectures built upon it.
Keywords:
APTx Neuron APTx Activation Function Activation Function Unified Neuron Deep Learning MNIST
1 Introduction
Custom activation functions such as Swish [13], Mish [9], and ELU [2] have demonstrated superior performance compared to the traditional ReLU [10] in various deep learning applications. These advancements reflect a broader trend toward more adaptive and expressive non-linearities.
Among recent innovations, the APTx activation function [5, 6] is notable for its parametric, trainable formulation, which can approximate multiple activation behaviors, including Swish and Mish, and can also resemble Tanh-like curves under certain parameter settings. As noted in recent surveys on trainable activation functions [1], such flexibility enables neural networks to better adapt to task-specific requirements during training. The mathematical formulation of the APTx activation function is given in Equation 1.
$y=(\alpha+\tanh(\beta x))\cdot\gamma x$ (1)
Here, $\alpha$, $\beta$, and $\gamma$ are learnable parameters. This formulation enables adaptation of the dynamic shape during training for the APTx activation function.
Furthermore, the APTx activation function can generate the SWISH($x,\rho$) activation function at parameters $\alpha=1,\beta=\rho/2,\text{and }\gamma=1/2$. Similarly, we can use the values $\alpha=1,\beta=1/2,\text{and }\gamma=1/2$ for the negative part, and $\alpha=1,\beta=1,\text{and }\gamma=1/2$ for the positive part if we want to closely approximate the MISH activation function.
In this work, we extend the APTx function from just an activation function to a full-fledged trainable neuron by adding a bias term $\delta$ and integrating the summation mechanism of a neuron. The result is a compact, expressive unit that handles both linear and non-linear transformations natively.
2 Background and Motivation
In standard feedforward neural networks, a neuron performs a two-step process: it computes a weighted sum of inputs followed by the application of a non-linear activation function [14]. Mathematically, this is often written as shown in Equation 2.
$y=\phi\left(\sum_{i=1}^{n}w_{i}x_{i}+b\right)$ (2)
where $w_{i}$ are the trainable weights, $b$ is the bias term, and $\phi(\cdot)$ is a non-linear activation function such as ReLU, Tanh, Swish, Mish, or APTx activation function.
While this modular design offers flexibility, it also imposes structural redundancy and increased memory overhead. The separation between linear and non-linear components requires additional layers and parameters, making it less efficient, especially in memory-constrained or latency-critical environments.
Moreover, traditional neurons rely on fixed activation functions that remain the same across the network or layer. This rigid formulation limits the network’s ability to adapt activation behavior dynamically based on the input distribution or training dynamics.
The APTx activation function [5, 6], due to its parametric nature, already offers adaptive behavior and a rich expressiveness to simulate or interpolate between several standard non-linearities. It introduces trainable parameters $\alpha$, $\beta$, and $\gamma$ into the activation itself, allowing it to learn optimal non-linear transformations during training. This results in improved flexibility and performance across tasks and architectures.
Extending this idea, we hypothesized that merging the activation and computation stages into a single unit, while preserving full trainability, can lead to more compact and powerful architectures. Instead of separating a weighted summation and a subsequent activation, we propose a unified formulation that naturally incorporates both. This approach eliminates the need for explicit activation layers, reduces parameter duplication, and potentially enhances representational efficiency.
This line of thinking led to the design of the APTx Neuron, described in the next section.
3 APTx Neuron
The APTx Neuron is a novel computational unit that unifies linear transformation and non-linear activation into a single, expressive formulation. Inspired by the parametric APTx activation function, this neuron architecture removes the strict separation between computation and activation, allowing both to be learned as a cohesive entity. It is designed to enhance representational flexibility while reducing architectural redundancy.
3.1 Mathematical Formulation
Traditionally, a neuron computes the output as shown in Equation 3.
$y=\phi\left(\sum_{i=1}^{n}w_{i}x_{i}+b\right)$ (3)
where $x_{i}$ are the inputs, $w_{i}$ are the weights, $b$ is the bias, and $\phi$ is an activation function such as ReLU, Swish, or Mish.
The APTx Neuron merges these components into a unified trainable expression, as shown in Equation 4.
$y=\sum_{i=1}^{n}\left[(\alpha_{i}+\tanh(\beta_{i}x_{i}))\cdot\gamma_{i}x_{i}\right]+\delta$ (4)
where:
  • $x_{i}$ is the $i$-th input feature,
  • $\alpha_{i}$, $\beta_{i}$, and $\gamma_{i}$ are trainable parameters for each input,
  • $\delta$ is a trainable scalar bias.
This equation allows the neuron to modulate each input through a learned, per-dimension non-linearity and scaling operation. The term $(\alpha_{i}+\tanh(\beta_{i}x_{i}))$ introduces adaptive gating, and $\gamma_{i}x_{i}$ provides multiplicative control.
3.2 Relationship to Traditional Neurons and Activations
A unique property of the APTx Neuron is its ability to represent multiple computational regimes depending on parameter values:
  • Linear Neuron Equivalence: If $\beta_{i}=0$, then $\tanh(\beta_{i}x_{i})=0$, and the equation reduces to:
    $y=\sum_{i=1}^{n}(\alpha_{i}\cdot\gamma_{i}x_{i})+\delta$ (5)
    If we further assume either $\gamma_{i}=1$ or $\alpha_{i}=1$, then it becomes:
    $y=\sum_{i=1}^{n}(\alpha_{i}\cdot x_{i})+\delta\quad\text{or}\quad y=\sum_{i=1}^{n}(\gamma_{i}\cdot x_{i})+\delta$ (6)
    which resembles the form of a conventional linear neuron with learnable weights and bias.
  • Pure Activation Behavior: When $\delta=0$ and input is passed as a fixed vector (e.g., $x=[x_{1},x_{2},...,x_{n}]$), the APTx Neuron computes a non-linear transformation akin to a composite APTx activation function:
    $y=\sum_{i=1}^{n}\left[(\alpha_{i}+\tanh(\beta_{i}x_{i}))\cdot\gamma_{i}x_{i}\right]$ (7)
  • Identity Function: When $\alpha_{i}=1/\gamma_{i}$, and $\beta_{i}=0$, the neuron becomes:
    $y=\sum_{i=1}^{n}x_{i}+\delta$ (8)
    again acting as a simple summing neuron with bias, matching standard behavior.
Because $\alpha_{i}$, $\beta_{i}$, $\gamma_{i}$, and $\delta$ are all trainable, the APTx Neuron can automatically adjust its role during training. It may act as a linear unit in some regions of the input space, and as a highly non-linear one in others.
3.3 Design Benefits
The unified formulation of APTx Neuron offers several practical and theoretical advantages:
  • Expressive Adaptivity: Each input dimension has its own dynamic non-linearity and transformation, enabling fine-grained learning.
  • Reduced Structural Complexity: APTx Neurons eliminate the need for separate activation layers in hidden layers by integrating non-linearity within the neuron itself.
  • Enhanced Generalization: The increased modeling freedom can help the APTx Neuron learn more compact representations without sacrificing performance.
  • Parameter Reusability: APTx Neurons are capable of mimicking multiple types of traditional neurons, eliminating the need for hand-picking activation functions.
3.4 Parameter Overhead and Efficiency
Each APTx Neuron introduces $13n+$ 1 trainable parameters for an input dimension $n$, compared to $n+1$ in a standard neuron. However, due to their richer expressiveness, fewer APTx Neurons or layers may be required to achieve comparable or better performance. This often leads to a favorable trade-off between parameter count and model accuracy.
Although the original formulation uses parameters $\alpha_{i}$, $\beta_{i}$, and $\gamma_{i}$ for each input dimension $x_{i}$ within a neuron, along with a shared parameter $\delta$, several parameter sharing strategies can be adopted to reduce model complexity:
  • Full sharing: All parameters are shared and trainable across input dimensions, i.e., $\alpha_{i}=\alpha$, $\beta_{i}=\beta$, $\gamma_{i}=\gamma$, and a shared, trainable $\delta$.
  • Partial sharing: For example, $\alpha_{i}=\alpha$ (shared and trainable), while $\beta_{i}$ and $\gamma_{i}$ remain input specific and trainable; $\delta$ remains shared and trainable.
  • Hybrid schemes: Certain parameters may be fixed. For instance, setting $\alpha_{i}=1$ (non-trainable), while keeping $\beta_{i}$, $\gamma_{i}$, and $\delta$ trainable. For a neuron with input dimension $n$, this reduces the total number of trainable parameters from $13n+$ 1 to $12n+$ 1.
These variants offer configurable trade-offs between parameter efficiency and expressive power, allowing flexible adaptation of the APTx Neuron to a variety of architectures and tasks.
Importantly, the formulation of the APTx Neuron preserves the universal approximation capability [4] of neural networks. By embedding trainable, input-wise non-linearities directly within each neuron, the APTx Neuron goes beyond traditional designs where approximation power depends on stacking fixed activations atop linear transformations. Each input dimension in the APTx Neuron can independently learn its own transformation behavior, both linear and non-linear, enabling more efficient, compact, and expressive modeling, even in shallow architectures.
Additionally, the APTx Neuron offers computational efficiency during training. As shown in Equation 4, the formulation of APTx Neuron relies on the $\tanh$ function, which is faster to compute than the sigmoid or softplus functions used in Swish and Mish. This makes the derivative easier to evaluate during backpropagation and reduces overall training overhead, providing a practical performance advantage without sacrificing expressiveness.
In general, the APTx Neuron offers a powerful, flexible, and theoretically grounded alternative to conventional neuron designs. Its ability to bridge between pure activation behavior, linear transformation, and hybrid modes makes it a strong candidate for building more efficient and expressive neural networks.
4 Architecture, Training, and Results
To evaluate the effectiveness of the proposed APTx Neuron, we implemented a custom fully connected feedforward neural network using APTx Neurons in PyTorch [11]. This section outlines the design, training pipeline, and performance metrics on the MNIST dataset [3].
4.1 Neural Network Design
The APTx Neuron-based feedforward neural network replaces conventional linear and activation layers with custom layers composed of multiple APTx Neurons. Each APTx Neuron unifies computation and non-linearity into a single trainable expression. In this configuration, all parameters $\alpha_{i}$, $\beta_{i}$, $\gamma_{i}$, and $\delta$ were trainable. The architecture used in our experiments is as follows:
  • Input: Flattened MNIST image of size $28\times 28=784$.
  • Layer 1: APTx Layer with 128 neurons.
  • Layer 2: APTx Layer with 64 neurons.
  • Layer 3: APTx Layer with 32 neurons.
  • Output Layer: Fully-connected linear layer with 10 output classes.
The total number of trainable parameters in the APTx Neuron-based feedforward neural network used in our experiments was 332,330. While the APTx layers unify activation and computation, a softmax function is applied to the final output layer to produce class probabilities in the classification task in the MNIST dataset [3].
The source code for the APTx Neuron-based architecture and MNIST [3] experiment, implemented in Python [12] using the PyTorch library [11], is available at our GitHub repository [7].
4.2 Training Details
  • Dataset: MNIST handwritten digit dataset.
  • Optimizer: Adam with initial learning rate $4\times 10^{-3}$.
  • Scheduler: StepLR with step size of 5 epochs and decay rate of 0.25.
  • Loss Function: CrossEntropyLoss.
  • Epochs: 20.
  • Batch Size: 64 for training, 1000 for testing.
All experiments were conducted on a single device (CPU or CUDA depending on availability), and training logs including losses and accuracy were recorded per epoch.
4.3 Performance Metrics
The APTx Neuron based model achieved a peak test accuracy of 96.69% at epoch 11. The training loss approached zero after epoch 8, and the training accuracy reached over 99.8% by epoch 20, indicating excellent capacity and convergence, as shown in Table 1. Visual analysis of the training and test loss values is shown in Figure 1, and the accuracy is shown in Figure 2.
Table 1: Performance of APTx Neuron-based feedforward neural network on MNIST.
Epoch Train Loss Test Loss Train Accuracy (%) Test Accuracy (%)
1 85.58 36.73 84.16 89.12
2 33.27 17.82 90.16 90.76
3 19.97 28.16 91.80 90.82
4 9.98 27.00 92.55 90.66
5 15.28 24.45 93.59 93.03
6 13.88 9.13 97.11 96.33
7 9.35 8.84 97.47 95.53
8 0.00 7.73 97.38 95.51
9 1.10 9.19 97.51 94.47
10 6.41 8.69 97.56 95.59
11 0.00 6.81 98.75 96.69
12 0.00 6.57 99.11 96.53
13 0.00 6.67 99.19 96.57
14 0.00 7.29 99.21 96.40
15 0.00 6.90 99.23 96.46
16 0.00 6.25 99.60 96.63
17 0.00 6.21 99.77 96.58
18 0.00 6.02 99.79 96.65
19 0.00 5.95 99.78 96.68
20 0.00 6.13 99.81 96.56
Figure 1: Visual analysis of train and test loss values.
Figure 2: Visual analysis of train and test accuracy values.
4.4 Insights
  • APTx Neuron-based feedforward neural network converges quickly. It surpasses 96% test accuracy within the first 6 epochs.
  • The near-zero training loss from epoch 8 onward suggests that the model has sufficient capacity to fully learn the MNIST data distribution.
  • Despite a higher number of parameters per neuron, the overall architecture remains compact and efficient.
  • The use of trainable non-linearities (APTx Neuron) enables superior representational power and dynamic learning behavior.
5 Toward Integration in CNNs and Transformers
The APTx Neuron, as introduced in this paper, is a general-purpose, unified computational unit that combines both activation and transformation into a single trainable expression. Although we have demonstrated its effectiveness within a fully connected feedforward neural network, the design is not limited to MLPs. As a fundamental building block, the APTx Neuron can be readily extended to modern architectures such as convolutional neural networks (CNNs) and transformers.
In CNNs [8], activation functions like ReLU or Swish are typically applied after convolutional filters. By replacing these fixed activations with APTx-style computation, it is possible to create APTx Convolutional Units that learn both spatial filtering and adaptive non-linearity in a unified manner, potentially improving spatial feature learning.
In transformer architectures [15], non-linear transformations are used in the feedforward sublayers and position-wise projections. Substituting those fixed activations with APTx Neurons may allow the model to dynamically adapt its internal representations and activation behavior based on the context or task, an especially valuable trait in attention-based models.
Thus, the APTx Neuron is not only a compact and expressive component for standard networks, but also a promising primitive for redefining core operations in CNNs, Transformers, and hybrid neural systems.
6 Conclusion
This work introduced the APTx Neuron, a unified, fully trainable neural unit that integrates linear transformation and non-linear activation into a single expression, extending the APTx activation function. By learning per-input parameters $\alpha_{i}$, $\beta_{i}$, $\gamma_{i}$, and $\delta$ for each input $x_{i}$ within a neuron, the APTx Neuron removes the need for separate activation layers and enables fine-grained input transformation. APTx Neuron generalizes traditional neurons and activations, offering greater representational power. Our MNIST experiments show that a fully connected APTx Neuron-based feedforward neural network achieves $96.69\%$ test accuracy in 20 epochs with approximately 332K trainable parameters, demonstrating rapid convergence and high efficiency. This design lays the groundwork for extending APTx Neurons to CNNs and transformers, paving the way for more compact and adaptive deep learning architectures.
{credits}
6.0.1 Competing Interest
The authors have no competing interests to declare that are relevant to the content of this article.
Funding Information
No institutional funding was received for this research.
Author Contribution
All authors have contributed equally to the conception, implementation, analysis, and writing of the manuscript.
Data Availability Statement
The source code and related experimental data used in this study are publicly available at: https://github.com/mr-ravin/APTxNeuron.
Research Involving Human and/or Animals
Not Applicable.
Informed Consent
Not Applicable.